OpenStudy (anonymous):

a 10-arce field produced 750 bushels of corn. at that rate, how much corn can be produced from a 14-arce field help? explain how?

OpenStudy (anonymous):

i believe this is a simple matter of isolating an unknown variable. so make 2 equations, one for each rate and make them equal to each other. 750/10 = x/14 (the 750/10 is the rate for the 10acre field and since we dont know the rate for the 14acre field, you use a variable x) x= (750/10)(14) (take 14 to the other side of the equation to isolate for x) x=1050